OpenStudy (anonymous):

-1/4x = 3/2 -1/4x = 3/2 @Mathematics

OpenStudy (anonymous):

-1/4x = 3/2, -4x/4x = 3(4x)/2, -1 = 12x/2, 6x = -1, x = -1/6